OpenStudy (anonymous):

Henry’s law constant for carbon dioxide at 38°C is 0.00228 mol L-1 atm-1. Calculate the pH of a solution of carbon dioxide at 38°C in equilibrium with the gas at a partial pressure of 3.20 atm

OpenStudy (anonymous):

ok so henry's law: p= k*c so you can now write: c=p/k = 3,2 atm / 0,00228 mol dm-3 atm-1 = 1403,5 mol dm-3 pH = -log c = -log 1403,5 = -3,15

OpenStudy (anonymous):

cool man thanks very much!!

OpenStudy (anonymous):

not a problem